# Kareem Amin  Alex Bie  Weiwei Kong  Alexey Kurakin Natalia Ponomareva  Umar Syed  Andreas Terzis  Sergei Vassilvitskii Google

### v1

@@ 1 Introduction
Private prediction methods do not modify the LLM parameters at all. Instead, they directly prompt the LLM with text from the source dataset, asking for similar text in response, and then perturb the LLM’s token distribution (i.e., its last layer) to ensure that each sampled token, and thus the entire generated response, is private. Since no training is required, private prediction methods can quickly generate synthetic data, typically producing some data within minutes instead of hours, which allows for rapid prototyping and iteration. However, unlike private fine-tuning, the guarantees of private prediction methods degrade with the volume of data that is generated. Consequently, existing private prediction methods have mostly been used in applications that require only small amounts of synthetic data [Tang et al., 2024], sharply limiting their practicality.
@@
  1. 1.
    Instead of protecting the privacy of the entire token distribution with Gaussian or Laplace noise, we leverage the uncertainty inherent in sampling to ensure privacy. We clip and aggregate token logits before standard softmax sampling — which is differentially private, since it can viewed as the exponential mechanism. Our approach induces much less distortion of the original token distributions to achieve the same level of privacy than prior work.
  2. 2.
    Previous work generated each token using a random subset of the input data, leveraging privacy amplification by subsampling in their analysis. This is computationally undesirable, as it requires repeated re-computation of the prefix for each decoding step, and limits scalability towards generating large synthetic corpora. Instead, we generate each synthetic example using a fixed disjoint subset of the input data, which yields substantial savings in privacy cost (by leveraging parallel composition) while allowing us to pay a linear amount of non-attention FLOPs, rather than quadratic, in terms of sequence length (via KV cache accelerated decoding).
  3. 3.
@@
Taken together, the combination of these algorithmic techniques leads to significant improvements over prior work. Informally, (1) and (2) above keep our inference closely aligned to standard (non-DP) inference.
In our experiments, we generate private synthetic versions of publicly available, benchmark machine learning datasets, and then use the synthetic datasets for downstream classification and extraction tasks. Owing to the increased quantity and quality of our synthetic data, we improve over an existing state-of-the-art private prediction method in terms of downstream accuracy. Furthermore, while prior work in this paradigm only generated a small (<10) number of examples we demonstrate the ability to generate thousands of training examples, enough for fine-tuning downstream models.
Finally, since synthetic data is intended for a wide variety of applications, we also evaluate data quality using a metric that is orthogonal to performance on a downstream classification task. Specifically, we generate synthetic versions of a publicly available dataset containing highly structured data records, each of which is encoded as a JSON object. Our results demonstrate that the sparse vector technique helps preserve data structure at high privacy levels.
@@ 2 Related work
Private fine-tuning is widely used for synthetic text generation. Yue et al. [2023] created private synthetic versions of text datasets by using DP-SGD [Abadi et al., 2016] to fine-tune an LLM on the sensitive data. Kurakin et al. [2024] showed that parameter efficient approaches to fine-tuning, such as LoRA [Hu et al., 2022], can improve the quality of the synthetic data, since reducing the number of parameters also reduces the amount of noise injected into the optimization procedure. Wu et al. [2024a] took a two-stage approach: First they fine-tuned an LLM on a public dataset that closely resembled the sensitive data (which was itself generated by an LLM using carefully designed prompts), and then completed the fine-tuning process by running DP-SGD on the sensitive dataset. Concurrent to this work, Tran and Xiong [2024] describe a private fine-tuning approach for generating synthetic tabular data that is formatting compliant.
@@
Private prediction [Dwork and Feldman, 2018] is an alternate approach to private machine learning that only guarantees the privacy of the predictions output by an ML model, and not the model itself. Private prediction has been applied to synthetic text generation by viewing each token sampled by an LLM as a ‘prediction’, and perturbing the LLM’s token distributions to ensure their privacy. Tang et al. [2024] added noise to several independent token distributions and averaged them, while Hong et al. [2024] selected the most popular token among the token distributions using the LimitedDomain mechanism [Durfee and Rogers, 2019]. These methods can avoid the time, compute, and access required to fine-tune an LLM with billions of parameters. However, a privacy loss is suffered for each token produced in this manner. As a result, previous work has only been able to generate a very small number of synthetic examples at reasonable privacy levels (fewer than 10). Other work has applied private prediction techniques to LLMs [Majmudar et al., 2022, Duan et al., 2023], including in combination with fine-tuning [Ginart et al., 2022, Flemings et al., 2024], but not for the purpose of synthetic text generation.
@@
Finally, another distinct set of approaches are private filtering methods. Private filtering methods operate directly on whole LLM responses and a large corpus of public data that does not require protection. Yu et al. [2024] and Xie et al. [2024] used the sensitive responses to privately select similar responses from the public dataset. Similarly, Wu et al. [2024b] aggregate response embeddings and select the public response that is closest in embedding space. [^1] Wu et al. [2024b] also proposed a non-filtering approach based on privately selecting common keywords among the sensitive data and using them to prompt an LLM. One limitation of filtering methods is that the menu of possible responses is constructed without signal from the new source dataset.
@@ 3 Method
Figure 1: Visualization of Algorithm 1 for a single token in a single batch.
Algorithm 1 describes our method for privately generating a dataset of synthetic examples $X$ from a dataset of sensitive prompts $D$. Each prompt in $D$ resembles the sample prompt given above. But instead of using a single prompt to generate a synthetic example, the algorithm uses a batch of the prompts to run several LLM inferences in parallel. Each synthetic example is generated one token at a time, with the average of the logit vectors across the inferences defining the distribution from which the next token is randomly selected. Before averaging, the logits are clipped and re-centered using the function
$\operatorname{clip}_{c}(\mathbf{z})_{i}=\max\left(-c,\mathbf{z}_{i}-\max_{j}{\{\mathbf{z}_{j}\}}+c\right)$ (1)
which maps each component of $\mathbf{z}$ into $[-c,c]$. Forcing each logit to lie in a bounded range is key to proving the privacy guarantee for our algorithm (see §4). While several functions can achieve this purpose, Eq. (1) has an additional desirable property: If the components of $\mathbf{z}$ can be shifted by a constant so that they all lie in the interval $[-c,c]$, then $\operatorname{clip}_{c}(\mathbf{z})$ is one such shift. This property is desirable because the distribution $\operatorname{softmax}(\mathbf{z})$ is invariant to any constant shift of $\mathbf{z}$. We also found that Eq. (1) performed better empirically than other functions we considered. For example, regular clipping to the range $[-c,c]$ without recentering requires twice as large $c$ to sample without distortion (see Appendix B).
Algorithm 1 Generate private synthetic examples using an LLM
1:Parameters: LLM logit function $\operatorname{logits}(\cdot)$, public prompt $\mathbf{p}_{\operatorname{public}}$, expected batch size $s$, maximum number of private tokens per batch $r$, clipping value $c$, noise level $\sigma$, distance function $d$, threshold $\theta$, public temperature $\tau_{\operatorname{public}}$, private temperature $\tau$
2:Input: Dataset of sensitive prompts $D$
@@
4:Let $\mathcal{S}$ be a partition of $D$ into disjoint batches
5:for each batch $S\in\mathcal{S}$ do
6:     $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
7:     $t\leftarrow 0$ private token counter
8:     while $t<r$ do
@@
10:         while $\mathbf{x}$ does not end with <eos> do
11:              $Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in S\}$
12:              $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$
13:              $\hat{d}\leftarrow d(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
14:              if $\hat{d}>\hat{\theta}$ then
15:                  $\bar{\mathbf{z}}\leftarrow\frac{1}{s}\sum_{\mathbf{z}\in Z}\operatorname{clip}_{c}(\mathbf{z})$
16:                  $x\sim\operatorname{softmax}(\bar{\mathbf{z}}/\tau)$
17:                  $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
@@
19:              else
20:                  $x\sim\operatorname{softmax}(\mathbf{z}_{\operatorname{public}}/\tau_{\operatorname{public}})$
21:              Append $x$ to $\mathbf{x}$
22:         Add $\mathbf{x}$ to $X$
23:Return $X$.
@@
Since the average logit vector is computed using a set of sensitive prompts, each token selected from a distribution determined by the average logit vector incurs a privacy cost. To minimize this cost, Algorithm 1 also has access to a non-sensitive public prompt, $\mathbf{p}_{\operatorname{public}}$, and uses this prompt to generate the next token whenever doing so does not significantly change the distribution from which the next token is drawn. The distance function used to make this determination is
$d(Z,\mathbf{z}_{\operatorname{public}})=\left\lVert\frac{1}{s}\sum_{\mathbf{z}\in Z}p_{\mathbf{z}}-p_{\mathbf{z}_{\operatorname{public}}}\right\rVert_{1}$ (2)
@@
where $p_{\mathbf{z}}=\operatorname{softmax}(\mathbf{z})$ is the token distribution corresponding to logit vector $\mathbf{z}$. When this distance is small, Algorithm 1 outputs a public token instead of a private token. The privacy guarantee for Algorithm 1 leverages the analysis of the sparse vector technique [Dwork et al., 2009], and shows that while privacy degrades with the number of private output tokens, it is independent of the number of public output tokens (see §4). Empirically, we observe that the fraction of output tokens that must be private in order to generate high-quality synthetic data can be only 20% for some datasets.
@@
Note that the first step of Algorithm 1 partitions the input dataset of sensitive prompts into disjoint batches. We do not prescribe a procedure for assigning prompts to batches in Algorithm 1 since many batching approaches are admissible as long as they satisfy a minor technical assumption required for the privacy analysis of Algorithm 1, which we explain in §4. While the batches are not required to be any particular size, the algorithm makes most efficient use of the data if each batch has size equal to the expected batch size $s$. And while prompts can be batched together (almost) arbitrarily, more tailored batching can lead to better synthetic data quality. For example, in the experiments in §5, where we generate synthetic versions of ML training datasets, each sensitive prompt contains a label. In those experiments we assign prompts with the same label to the same batch.
Relationship to prior work.
@@
Two major features of Algorithm 1 are that it leverages the inherent randomness of token sampling to guarantee privacy, and that it further reduces privacy cost by using public data to generate a portion of the synthetic data. Some prior work also incorporated these algorithmic ideas, but with key differences. Instead of clipping logits to ensure that the token sampling is private, Majmudar et al. [2022] mixed each sensitive token distribution with the uniform distribution. This approach induced a dependence on the vocabulary size in their privacy guarantee, and since LLM vocabularies are typically very large, the resulting privacy guarantee was quite weak: Majmudar et al. [2022] noted that setting the differential privacy parameter $\varepsilon$ (see Definition 1) lower than 50 produced synthetic data that was “unusable”. Flemings et al. [2024] guaranteed the privacy of token sampling by mixing each sensitive token distribution with a public token distribution, but their approach was based on aggregating a set of fine-tuned models, not a set of prompts. Neither Majmudar et al. [2022] nor Flemings et al. [2024] had a goal of generating synthetic data.
@@
Tang et al. [2024] found that limiting the token vocabulary to a fixed set of the most popular 100 public tokens caused their synthetic data generation algorithm to exhibit greater stability. However, if the sensitive data contains many tokens that are rare in public data, their approach cannot produce synthetic data that is very similar to the sensitive data. By contrast, our approach compares public and private token distributions on-the-fly, and determines which one to use for sampling the next token by balancing a trade-off between privacy and quality. Also, Tang et al. [2024] used a different random subset of prompts to generate each token, and left as an open problem how to use a single subset to generate every token in a synthetic example. Our algorithm resolves this open problem, and consequently yields both improved privacy and greater computational efficiency (see §6).

### v2

# Private prediction for large-scale synthetic text generation

@@ Abstract
Private prediction for large-scale synthetic text generationthanks: Authors ordered alphabetically. Author contributions are listed at the end.
Kareem Amin Alex Bie Weiwei Kong Alexey Kurakin Natalia Ponomareva Umar Syed Andreas Terzis Sergei Vassilvitskii Google {kamin,alexbie,weiweikong,kurakin,nponomareva,usyed,aterzis,sergeiv}@google.com
@@ 1 Introduction
Private prediction methods do not modify the LLM parameters at all. Instead, they directly prompt the LLM with text from the source dataset, asking for similar text in response, and then perturb the LLM’s token distribution (i.e., its last layer) to ensure that each sampled token, and thus the entire generated response, is private. Since no training is required, private prediction methods can quickly generate synthetic data, typically producing some data within minutes instead of hours, which allows for rapid prototyping and iteration. However, unlike private fine-tuning, the guarantees of private prediction methods degrade with the volume of data that is generated. Consequently, existing private prediction methods have mostly been used in applications that require only small amounts of synthetic data (Tang et al., 2024), sharply limiting their practicality.
Figure 1: Algorithm 1, visualized. An LLM receives a batch of prompts, each instructing to generate text similar to a piece of sensitive text. Synthetic text is generated token by token, by running inference on the batch in parallel. In each step, the logit vectors produced downstream of sensitive text are aggregated and sampled from with differential privacy. Every token sampled in such way incurs a privacy cost, motivating us to include an auxillary public prompt and sample from its logits when similar to the sensitive logits.
1.1 Our contributions
@@
1. Better private token selection.
Instead of protecting the privacy of the entire token distribution with Gaussian or Laplace noise, we leverage the uncertainty inherent in sampling to ensure privacy. We clip and aggregate token logits before standard softmax sampling — which is differentially private, since it can be viewed as the exponential mechanism. Our approach induces much less distortion of the original token distributions to achieve the same level of privacy than prior work.
2. Avoiding prefix re-sampling.
Prior work generated each token using a random subset of the input data, leveraging privacy amplification by subsampling in their analysis. This is computationally undesirable, as it requires repeated re-computation of the prefix for each decoding step, thus limiting scalability towards generating large synthetic corpora. Indeed, prior work describes this re-sampling as the “main weakness” of the approach Tang et al. (2024). To resolve the problem, we instead generate each synthetic example using a fixed disjoint subset of the input data, which yields substantial savings in privacy cost – via parallel composition – while allowing us to pay linear instead of quadratic non-attention FLOPs in terms of sequence length via KV cache accelerated decoding.
3. Leveraging public predictions.
@@
Taken together, the combination of these algorithmic techniques leads to significant improvements over prior work. Roughly speaking, (1) and (2) above keep our inference closely aligned to standard (non-DP) inference.
In our experiments, we generate private synthetic versions of publicly available, benchmark machine learning datasets, and then use the synthetic datasets for downstream classification and extraction tasks. Owing to the increased quantity and quality of our synthetic data, we improve over an existing state-of-the-art private prediction method in terms of downstream accuracy. Furthermore, while prior work in this paradigm only generated a small (<10) number of examples, we demonstrate the ability to generate thousands of training examples, enough for fine-tuning downstream models.
Finally, since synthetic data is intended for a wide variety of applications, we also evaluate data quality using a metric that is orthogonal to performance on downstream tasks. Specifically, we generate synthetic versions of a publicly available dataset containing highly structured data records, each of which is encoded as a JSON object. Our results demonstrate that the sparse vector technique helps preserve data structure at high privacy levels.
@@ 2 Related work
Private fine-tuning is widely used for synthetic text generation. Yue et al. (2023) created private synthetic versions of text datasets by using DP-SGD (Abadi et al., 2016) to fine-tune an LLM on the sensitive data. Kurakin et al. (2024) showed that parameter efficient approaches to fine-tuning, such as LoRA (Hu et al., 2022) can improve the quality of the synthetic data. Wu et al. (2024a) took a two-stage approach: First they fine-tuned an LLM on a public dataset that closely resembled the sensitive data (which was itself generated by an LLM using carefully designed prompts); then they completed the fine-tuning process by running DP-SGD on the sensitive dataset. Concurrent to the present work, Tran and Xiong (2024) describe a private fine-tuning approach for generating synthetic tabular data that is formatting compliant.
Private prediction (Dwork and Feldman, 2018) is an alternate approach to private machine learning that only guarantees the privacy of the predictions output by an ML model, and not the model itself. The predominant way this is realized is via subsample-and-aggregate (Nissim et al., 2007): First sensitive data is split into disjoint partitions; then non-private predictions are made from each partition and privately aggregated. PATE (Papernot et al., 2017, 2018) employs this approach to get answers to a limited set of image classification queries, which are then used to train a student model that can be queried indefinitely.
@@
Private prediction has been applied to synthetic text generation by viewing each token sampled by an LLM as a ‘prediction’, and perturbing the LLM’s token distributions to ensure their privacy. Tang et al. (2024) added noise to several independent token distributions and averaged them, while Hong et al. (2024) selected the most popular token among the token distributions using the LimitedDomain mechanism (Durfee and Rogers, 2019). These methods can avoid the time, compute, and access required to fine-tune an LLM with billions of parameters. However, a privacy loss is suffered for each token produced in this manner. As a result, previous work has only been able to generate a very small number of synthetic examples at reasonable privacy levels (fewer than 10). Other work has applied private prediction techniques to LLMs (Majmudar et al., 2022; Duan et al., 2023), including in combination with fine-tuning (Ginart et al., 2022; Flemings et al., 2024), but not for the purpose of synthetic text generation.
@@
Finally, another distinct set of approaches are private filtering methods. Private filtering methods operate directly on whole LLM responses and a large corpus of public data that does not require protection. Yu et al. (2024) and Xie et al. (2024) used the sensitive responses to privately select similar responses from the public dataset. Similarly, Wu et al. (2024b) aggregate response embeddings and select the public response that is closest in embedding space. [^1] Wu et al. (2024b) also proposes a non-filtering approach based on privately selecting common keywords among the sensitive data and using them to prompt an LLM. One limitation of filtering methods is that the menu of possible responses is constructed without signal from the new source dataset.
@@ 3.2 Our algorithm
Algorithm 1 describes our method for privately generating a dataset of synthetic examples $X$ from a dataset of sensitive prompts $D$. Each prompt in $D$ resembles the sample prompt given above. But instead of using a single prompt to generate a synthetic example, Algorithm 1 takes a batch of prompts $S$ and runs LLM inference in parallel on each prompt. A synthetic example is generated one token at a time, with the average of the logit vectors across the batch defining the distribution from which the next token is randomly selected. Before averaging, a logit vector $\mathbf{z}$ is clipped and re-centered using the function
$\operatorname{clip}_{c}(\mathbf{z})_{i}=\max\{-c,\mathbf{z}_{i}-\max_{j}{\{\mathbf{z}_{j}\}}+c\}$ (1)
which maps each component $i$ of $\mathbf{z}$ into the target clipping range $[-c,c]$. Forcing each logit to lie in a bounded range is key to proving the privacy guarantee for our algorithm (see §4). While several functions can achieve this purpose, Eq. (1) has an additional desirable property: If the components of $\mathbf{z}$ can be shifted by a constant so that they all lie in the interval $[-c,c]$, then $\operatorname{clip}_{c}(\mathbf{z})$ is one such shift. This property is desirable because the distribution $\operatorname{softmax}(\mathbf{z})$ is invariant to any constant shift of $\mathbf{z}$. Empirically, we found that Eq. (1) performed better than other functions considered. For example, regular clipping to the range $[-c,c]$ without recentering requires twice as large $c$ to sample without distortion (see §B.1).
Algorithm 1 Generate private synthetic examples
1:Parameters: LLM $\operatorname{logits}(\cdot)$, public prompt $\mathbf{p}_{\operatorname{public}}$, expected batch size $s$, private tokens to sample $r$. Sampling: clipping bound $c$, temperature $\tau$, public temperature $\tau_{\operatorname{public}}$. Public optionality: public/private distribution distance $\mathrm{d}(\cdot,\cdot)$, threshold $\theta$, noise level $\sigma$.
2:Input: Dataset of sensitive prompts $D$; each prompt contains a sensitive example
@@
4:$X\leftarrow\emptyset$
5:Let $\mathcal{S}$ be a partition of $D$ into disjoint batches
6:for each batch $S\in\mathcal{S}$ do
7:     $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$# Init noisy threshold
8:     $t\leftarrow 0$# Private token counter
9:     while $t<r$ do
@@
11:         while $\mathbf{x}$ does not end with <eos> do
12:              $Z\leftarrow\{\operatorname{logits}(\mathbf{p}\mathbf{x}):\mathbf{p}\in S\}$
13:              $\mathbf{z}_{\operatorname{public}}\leftarrow\operatorname{logits}(\mathbf{p}_{\operatorname{public}}\mathbf{x})$     # Check if pub/priv distributions are far
14:              $\hat{d}\leftarrow\mathrm{d}(Z,\mathbf{z}_{\operatorname{public}})+\textrm{Laplace}(2\sigma)$
15:              if $\hat{d}\geq\hat{\theta}$ then# Sample priv token
16:                  $\bar{\mathbf{z}}\leftarrow\frac{1}{s}\sum_{\mathbf{z}\in Z}\operatorname{clip}_{c}(\mathbf{z})$
17:                  $x\sim\operatorname{softmax}(\bar{\mathbf{z}}/\tau)$
@@
19:                  $\hat{\theta}\leftarrow\theta+\textrm{Laplace}(\sigma)$
20:              else# Sample pub token
21:                  $x\sim\operatorname{softmax}(\mathbf{z}_{\operatorname{public}}/\tau_{\operatorname{public}})$
22:              Append $x$ to $\mathbf{x}$
23:         $X\leftarrow X\cup\{\mathbf{x}\}$
24:return $X$
@@
Since the average logit vector is computed using sensitive prompts, each token selected from a distribution determined by the average logit vector incurs a privacy cost. To minimize this cost, Algorithm 1 also has access to a non-sensitive public prompt, $\mathbf{p}_{\operatorname{public}}$, and uses this prompt to generate the next token whenever doing so does not significantly change the distribution from which the next token is drawn. The distance function used to make this determination is
$\mathrm{d}(Z,\mathbf{z}_{\operatorname{public}})=\|\frac{1}{s}\sum_{\mathbf{z}\in Z}p_{\mathbf{z}}-p_{\mathbf{z}_{\operatorname{public}}}\|_{1},$ (2)
@@
where $p_{\mathbf{z}}:=\operatorname{softmax}(\mathbf{z})$, $Z$ are the logit vectors computed for each sensitive prompt in $S$, $\mathbf{z}_{\operatorname{public}}$ is the logit vector computed using $\mathbf{p}_{\operatorname{public}}$, and $s$ is the expected batch size. When this distance is small, Algorithm 1 outputs a public token instead of a private token. The privacy guarantee for Algorithm 1 leverages the analysis of the sparse vector technique (Dwork et al., 2009), and shows that while privacy degrades with the number of private output tokens, it is independent of the number of public output tokens (see §4). Empirically, we observe that the fraction of output tokens that must be private in order to generate high-quality synthetic data can be as low as 20% for highly structured datasets.
@@
Note that the first step of Algorithm 1 partitions the input dataset of sensitive prompts into disjoint batches. We do not prescribe a procedure for assigning prompts to batches in Algorithm 1 since many batching approaches are admissible as long as they satisfy a minor technical assumption required for the privacy analysis of Algorithm 1, which we explain in §4. While the batches are not required to be any particular size, the algorithm runs faster if each batch has size equal to the expected batch size $s$. And while prompts can be batched together (almost) arbitrarily, more tailored batching can lead to better synthetic data quality. For example, in the experiments in §5, where we generate synthetic versions of ML training datasets, each sensitive prompt contains a label. In those experiments we assign prompts with the same label to the same batch.
3.3 Comparison to prior algorithms
@@
Two major features of Algorithm 1 are that it leverages the inherent randomness of token sampling to guarantee privacy, and that it further reduces privacy cost by using public data to generate a portion of the synthetic data. Some prior work also incorporated these algorithmic ideas, but with key differences. Instead of clipping logits to ensure that the token sampling is private, Majmudar et al. (2022) mixed each sensitive token distribution with the uniform distribution. This approach induced a dependence on the vocabulary size in their privacy guarantee, and since LLM vocabularies are typically very large, the resulting privacy guarantee was quite weak: Majmudar et al. (2022) noted that setting the differential privacy parameter $\varepsilon$ lower than 50 produced synthetic data that was “unusable”. Flemings et al. (2024) guaranteed the privacy of token sampling by mixing each sensitive token distribution with a public token distribution, but their approach was based on aggregating a set of fine-tuned models, not a set of prompts. Neither Majmudar et al. (2022) nor Flemings et al. (2024) aim to generate synthetic data.
@@
Tang et al. (2024) found that limiting the token vocabulary to a fixed set of the most popular 100 public tokens caused their synthetic data generation algorithm to exhibit greater stability. However, if the sensitive data contains many tokens that are rare in public data, their approach cannot produce synthetic data that is very similar to the sensitive data. By contrast, our approach compares public and private token distributions on-the-fly, and determines which one to use for sampling the next token by balancing a trade-off between privacy and quality. Also, Tang et al. (2024) used a different random subset of prompts to generate each token, and left as an open problem how to use a single subset to generate every token in a synthetic example. Our algorithm resolves this open problem, and consequently yields both improved privacy and greater computational efficiency (see §6).